40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 1527     Entry time: Thu Sep 25 16:50:09 2014
Author: Evan 
Type: DailyProgress 
Category: FSS 
Subject: Should probably move south BB EOM 

We've noticed for a while now that we cannot turn up the gain on the south TTFSS as high as on the north TTFSS, despite having similar optical power levels, similar mode-matching, etc. (See the OLTFs in ctn:1504.) The north gain can be set to 900/900 on the common/fast knobpots, but on south it's more like 600/600.

Because the BB EOM is placed before the PMC, I suspect the cavity pole of the PMC (1.8 MHz, measured in elog: in 2010) is giving us extra phase which prevents us from turning the loop gain up higher. Indeed, when I remove the PMC from the south optical path (and realign into the south cavity) I find I can turn the south TTFSS knobpots up to 800/800. A new OLTF is probably in order.

The easy thing to do for now is to leave the PMC out. The better thing is probably to move the BB EOM to come after the PMC. Since there's no room, this probably means putting the BB EOM where the resonant EOM currently is, put the resonant EOM where the EOAM currently is, and then put the EOAM elsewhere. The EOAM could just as well come before the PMC, since we're only attempting intensity stabilization well below the PMC cavity pole.

ELOG V3.1.3-